# Symbolic-AI-Fusion Deep Learning (SAIF-DL): Encoding Knowledge into Training with Answer Set Programming Loss Penalties by a Novel Loss Function Approach

### v1

@@ 1 Introduction
In summary, our contribution lies in presenting a scalable and flexible hybrid approach that seamlessly integrates domain expert knowledge into the training of DL models. This methodology is applicable to both regression and classification tasks and can be generalized across diverse fields, particularly those requiring strict adherence to domain constraints.
@@ 2 Related Work
Purely data-driven modelling using neural networks is agnostic with respect to the laws and rules underlying the data and, by design, primarily meant for interpolation, with a comparably restricted potential for transfer learning. As a consequence, purely data-driven modelling requires a large volume of input data in order to sufficiently cover the domain within which interpolation is to be facilitated. As another consequence, it may fail to correctly reproduce facts or internal consistency relationships that appear trivial to the end user, creating embarrassing counterexamples. Neural-symbolic approaches seek to bridge these limitations by combining the adaptive learning capabilities of neural networks with the logical, rule-based reasoning of symbolic AI. This integration allows AI systems to tackle a broader range of tasks, encompassing both low-level perception (e.g., image and speech recognition) and high-level reasoning (e.g., inference and abstract concept manipulation) [4]. For instance, neuro-symbolic systems have shown promising results in visual question answering (VQA), where systems must interpret visual data and respond to natural language queries, blending perceptual and symbolic reasoning to achieve accurate results [8].
@@
Various symbolic AI models have been developed, each with unique strengths in knowledge representation and reasoning. For instance, logic programming using Prolog [14] supports problem-solving and formal reasoning by allowing the definition of facts, rules, and queries. Similarly, ASP emphasizes nonmonotonic reasoning for dynamic problem-solving in areas such as planning and constraint satisfaction [1]. One significant area of development within neural-symbolic AI is concept learning, which merges the pattern recognition strengths of neural networks with the abstract representation capabilities of symbolic reasoning. This combination is particularly beneficial for compositional reasoning tasks, where the goal is to understand relationships between different concepts. Notable systems, such as the Neuro-Symbolic Concept Learner (NS-CL), effectively integrate scene understanding with symbolic reasoning to perform tasks like concept learning and relational reasoning, thereby enabling AI models to generalize knowledge to new contexts [17]. A significant advancement in this domain is the advent of deep probabilistic programming languages (DPPLs) [16], which merge neural predicates into a probabilistic logic programming framework. This fusion allows for the flexibility of deep learning while incorporating the structured reasoning strengths of probabilistic inference, facilitating both symbolic and subsymbolic reasoning. NeurASP enhances traditional Answer Set Programs by utilizing neural networks to manage uncertainties in symbolic reasoning, interpreting neural outputs as probabilistic distributions for atomic facts. This method has been applied in areas such as probabilistic knowledge representation and decision-making systems [18]. In addition, physics-informed neural networks (PINNs) [5] are a class of neural networks that integrate fundamental physical laws into the learning process of deep learning models. Instead of relying solely on data to train the model, PINNs incorporate physical principles described by partial differential equations (PDEs) or other mathematical formulations directly into the network’s architecture or loss function. This approach ensures that the model’s predictions are consistent with known physics, enhancing both accuracy and generalizability.
@@ 4.1.1 Example of a Differentiable Penalty Function
Consider the same example in the context of battery manufacturing:

### v2

@@ Abstract
Keywords:
Answer set programming neural-symbolic learning XAIR models
@@ 2 Related Work
Purely data-driven modelling using neural networks is agnostic with respect to the laws and rules underlying the data and, by design, primarily meant for interpolation, with a comparably restricted potential for transfer learning. As a consequence, purely data-driven modelling requires a large volume of input data in order to sufficiently cover the domain within which interpolation is to be facilitated. As another consequence, it may fail to correctly reproduce facts or internal consistency relationships that appear trivial to the end user, creating embarrassing counterexamples.
Neural-symbolic approaches seek to bridge these limitations by combining the adaptive learning capabilities of neural networks with the logical, rule-based reasoning of symbolic AI. This integration allows AI systems to tackle a broader range of tasks, encompassing both low-level perception (e.g., image and speech recognition) and high-level reasoning (e.g., inference and abstract concept manipulation) [4]. For instance, neuro-symbolic systems have shown promising results in visual question answering (VQA), where systems must interpret visual data and respond to natural language queries, blending perceptual and symbolic reasoning to achieve accurate results [8]. Additionaly, such approaches perform well with limited training data due to the guidance provided by symbolic reasoners [1].
@@
Various symbolic AI models have been developed, each with unique strengths in knowledge representation and reasoning. For instance, logic programming using Prolog [14] supports problem-solving and formal reasoning by allowing the definition of facts, rules, and queries. Similarly, ASP emphasizes nonmonotonic reasoning for dynamic problem-solving in areas such as planning and constraint satisfaction [1]. One significant area of development within neural-symbolic AI is concept learning, which merges the pattern recognition strengths of neural networks with the abstract representation capabilities of symbolic reasoning. This combination is particularly beneficial for compositional reasoning tasks, where the goal is to understand relationships between different concepts. Notable systems, such as the Neuro-Symbolic Concept Learner (NS-CL), effectively integrate scene understanding with symbolic reasoning to perform tasks like concept learning and relational reasoning, thereby enabling AI models to generalize knowledge to new contexts [17]. A significant advancement in this domain is the advent of deep probabilistic programming languages (DPPLs) [16], which merge neural predicates into a probabilistic logic programming framework. This fusion allows for the flexibility of deep learning while incorporating the structured reasoning strengths of probabilistic inference, facilitating both symbolic and subsymbolic reasoning. NeurASP [18] enhances traditional Answer Set Programs by utilizing neural networks to manage uncertainties in symbolic reasoning, interpreting neural outputs as probabilistic distributions for atomic facts. This method has been applied in areas such as probabilistic knowledge representation and decision-making systems. In addition, physics-informed neural networks (PINNs) [5] are a class of neural networks that integrate fundamental physical laws into the learning process of deep learning models. Instead of relying solely on data to train the model, PINNs incorporate physical principles described by partial differential equations (PDEs) or other mathematical formulations directly into the network’s architecture or loss function. This approach ensures that the model’s predictions are consistent with known physics, enhancing both accuracy and generalizability.
@@ 4.1.1 Example of a Differentiable Penalty Function
Consider the example in the context of battery manufacturing:
@@ 5 Proof of Concept Experiment
Figure 2: Comparison of Training and Validation Loss Curves for Models With and Without ASP Penalty
Figure 2 illustrates the training and validation loss curves for models trained with and without the ASP penalty. In the left plot, without ASP, the training loss decreases steadily, but the validation loss flattens earlier, suggesting potential overfitting and moderate generalization. In contrast, the right plot, with ASP, shows closely aligned training and validation losses, highlighting reduced overfitting and enhanced generalization.
@@ 7 Conclusion
{credits}
7.0.1 Acknowledgements
Funding is acknowledged from the European Union’s Horizon Europe research and innovation programme under grant agreements no. 101137725 (BatCAT) and 101138510 (DigiPass CSA).
[…]
The authors have submitted a grant proposal in which the presented architecture plays a central role.